A parallelogram has a perimeter of 48 kilometers. Two of the sides are 9 kilometers long. What is the length of the other two sides?
Found 2 solutions by ewatrrr, harpazo:
Answer by ewatrrr(24785)

P = 48km
48-2•9 = 30
30/2 = 15km each are the other two sides

Answer by harpazo(655)

P = 2l + 2w
48 = 2(9) + 2w
48 = 18 + 2w
48-18 = 2w
30 = 2w
30/2 = w
15 = w
The other two sides are 15 kilometers.